OpenStudy (anonymous):

Detrmine the (dy/dx) for y^5-3x^2=x; Can anyone help me out with this one?

OpenStudy (anonymous):

5y^4(dx/dy) - 6x = 1 1 + 6x = 5y^4(dx/dy) (dx/dy) = ( 1 + 6x ) / ( 5y^ 4)

OpenStudy (vijay):

dy/ dx = 5y^4 / 1+6x